| Operation Summary for April 20, 2012 - April 23, 2012 |
|
NuMI reported receiving 7.92E18 protons on target for the period from 4/16/12 to 4/23/12. Friday April 20 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), with Muon Ring personnel conducting a study, and with the NuMI, MiniBooNE, SeaQuest, and Meson FTBF T-1015 experiments taking beam. During the day shift, Operators had the Neutrino Area and SeaQuest ready for access by 8:28 AM. The NuMI beam permit tripped off for the first of two times in about 3 hours at 8:44 AM due to a power supply (HV101) tripping off; everything reset. SeaQuest came out of access and resumed taking beam around 9:45 AM. At 9:56 AM, Operations turned off the Booster beam for about 20 minutes to allow RF technicians to bring a Booster RF station (BRF10) back online after repairs. A MI person began a MI-52 Septa beam scan around 11 AM. Operations established beam to MTA at 12:13 PM. T-1015 made access during this shift. During the evening shift, at 4:29 PM, Operations received an Operating Note to allow the Muon Ring to take 120 GeV beam to its target. Operations disabled beam to MTA at 6:05 PM. T-1015 and SeaQuest made accesses during this shift. Saturday April 21 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1015. During the midnight shift, at 12:20 AM, BRF11 tripped off for the first of four times in about three hours; it reset. At 12:28 AM, Operations completed a 750 KeV area beam line tune, but Operators only kept the changes from the Linac steer. Operations completed a Booster injection tune at 3:26 AM. Operations adjusted the Booster transition times at 4:30 AM. The NuMI beam permit tripped off at 4:46 AM due to a quadrupole power supply (E:QF107) drifting high; everything reset. During the day shift, at 8:10 AM, the NuMI beam permit tripped off due to E:QF115 drifting low; everything reset. At 8:25 AM, BRF11 tripped off on a ground fault indication for the first of five times in six hours; it reset. Operations established beam to MTA at 9:30 AM. Operations completed a Booster injection tune at 11:43 AM. At 12:23 PM, the NuMI beam permit tripped off due to E:HV101 tripping off; everything reset. Operations disabled beam to MTA at 3:49 PM. T-1015 and SeaQuest made accesses during this shift. During the evening shift, BRF11 tripped off at 8:55 PM; it reset. At 10:44 PM, the NuMI beam permit tripped off due to a bad beam position; it reset. At 11:40 PM, the NuMI beam permit tripped off when E:V118 and E:HV101 tripped off; everything reset. Operations reported at 11:48 PM that they had completed a 750 KeV tune. T-1015 and SeaQuest made accesses during this shift. Sunday April 22 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1015. During the midnight shift, Operations reported adjusting the Linac low-energy RF intertank phases at 3:12 AM to improve beam efficiencies. Operations completed a Booster injection tune at 3:12 AM. T-1015 made an access during this shift. During the day shift, at 8:45 AM, BRF11 tripped off three times in six minutes and would not reset after its last trip; Operators reset the station locally. Operations established beam to MTA at 10:18 AM. The MiniBooNE beam tripped off at 12:19 PM due to an airflow alarm (E:12AAIR); everything reset. Operations disabled the MTA beam at 3:05 PM. T-1015 and SeaQuest made accesses during this shift. During the evening shift, Operations reported multiple Booster RF trips on stations 11, 15, and 19 during this shift. Monday April 23 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1015. During the midnight shift, Operations turned off beam to MiniBooNE around 1:32 AM; this marked the end of the experiment’s beam for this year. At 5:58 AM, the NuMI beam permit tripped off due to HV101; everything reset. At 6:35 AM, the NuMI beam permit tripped off due to V118, everything reset. Operations Coordinator (OpsCo) Plans for Today The OpsCo agrees with the plans of Booster, Muon Ring, and MI personnel to conduct studies. SeaQuest continues to commission their beam line. MiniBooNE began its shutdown early. |
|
More Information Comments and Suggestions |
| last modified 04/23/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |